If 9 - 6x = 45 find the value of x - 4
Answer by robertb(4012)

If 9 - 6x = 45, then
6x - 9 = -45
==> 6x - 9 - 15 = -45 - 15
<==> 6x - 24 = -60
==> 6(x-4) = -60
==> x -4 = -10.